OpenStudy (anonymous):

simplify the question cube root of and 54 square rooted

OpenStudy (anonymous):

I got 9 and 6 for the square root of 54 so what would I do with the 3 cube root

OpenStudy (anonymous):

\[\sqrt[3]{\sqrt{54}} = \sqrt[3]{\sqrt{9*6}} = \sqrt[3]{3\sqrt{6}}\]

OpenStudy (anonymous):

okay now I get it thanks

OpenStudy (anonymous):

you could type that as: \[3^{1/3}6^{1/6}\]

OpenStudy (anonymous):

ok